what must be added to x^2 - x + 4 to male it equal to (x+2)^2

Answer by mananth(16946)

x^2 - x + 4 to make it equal to (x+2)^2
x^2 - x + 4 to make it equal to (x^2+4x+4)
X^2-x+4 +A = x^2+4x+4
A = x^2+4x+4 -x^2+x-4
A= 5x